# After Retrieval, Before Generation: Enhancing the␣Trustworthiness of Large Language Models in RAG

### v1

@@ 5.2 Experiment Results
Ablation Study. We perform an ablation study to evaluate the significance of Allocator and Decision Tree as shown in Table 2. (1) w/o Allocator: We remove the Allocator module, setting both $r_{p}$ and $g_{p}$ to 100% in BRIDGE. This eliminates soft bias guidance for knowledge collection and evaluation. The results show performance degradation in knowledge integration RAG scenarios (FA, FI, FE) within TRD compared to BRIDGE’s optimal performance.
@@
(2) w/o Decision Tree (ICL) & w/o Decision Tree (GRPO): Disabling the Maximum Soft-bias Decision Tree, we directly provide the LLM with knowledge and matching scores for final response generation. While this improved performance in some knowledge-integration scenarios, it caused significant performance drops in RA scenarios. The ablation study demonstrates that both modules play a critical role in enabling the LLM to maintain balanced response capabilities across all four RAG scenarios.
Allocator Performance. To further investigate the Allocator’s effectiveness, we examine its allocation behavior in BRIDGEGRPO mode across different RAG scenarios. As shown in Figure 5, the Allocator demonstrates accurate classification capability. The predominant allocation patterns of [80%-20%] and [20%-80%] indicate that the model can establish a correct soft bias direction for different RAG scenarios. This pronounced bias allocation significantly enhances the decision-making process within the decision tree, contributing to a reliable response.
Efficiency Analysis. In our experiments, the reflection mechanism is activated in approximately 10% of all cases. As shown in Figure 6, within these activated instances, the model’s problem-solving accuracy positively correlates with the number of reflection iterations. However, as a plug-and-play component, we carefully considered the efficiency trade-offs. Under ideal conditions, BRIDGE can resolve questions with 5 LLM API calls. When accounting for reflection operations, the average number of API calls per query reaches 5.89 on the TRD dataset - a number comparable to other prompt-based methods. Importantly, this modest increase in computational overhead enables BRIDGE to maintain balanced performance across diverse scenarios. We discuss the efficiency of the Reflection module in Appendix C.
Figure 6: Acc under different reflection iterations.
@@ 6 Conclusion
This paper addresses the critical challenge of balancing internal and external knowledge in retrieval-augmented generation (RAG) systems. First, we construct TRD to enable comprehensive evaluation. Then, we propose BRIDGE, a unified framework that introduces soft bias to dynamically weight knowledge sources, an Allocator module to guide retrieval/generation, and a Maximum Soft-bias Decision Tree to select optimal response strategies. Experiments demonstrate BRIDGE’s superiority over baselines, achieving balanced performance across all cases while significantly improving refusal rates in adversarial settings. Our work provides an effective and balanced approach to enhance the trustworthiness and robustness of RAG systems in real-world applications.
@@ Appendix A Trustworthiness Response Dataset
A.1 Data Generation for Different RAG Scenarios.
@@ A.1.1 FA Data
The NQ training set consists of 104,072 data samples. Following the method described in longpre2021entity , we perform entity matching based on popularity using the short-answer entities to generate two incorrect answers of the same type as the correct entity. We filter out invalid questions with empty long answers and samples where fewer than two corresponding entities could be retrieved, resulting in 76,409 valid QA pairs. For external knowledge, we use the provided long answers. For internal knowledge generation, referring to the prompt template of ConflictBank su2024conflictbank , we employ GPT-4-turbo with the following prompt template:
Prompt for Internal Knowledge Generation Based on the question and answer, create a piece of evidence that answers the question. 1. You cannot answer this question yourself. You must create evidence based on the answer I give. 2. This piece of evidence should be informative and well-structured. 3. The evidence should be presented in the tone of a Wikipedia entry. Question: {question} Answer: {correct_answer} Evidence:
@@ A.1.2 FI Data
The NQ validation set contains 12,837 samples. Following a similar method to FA, we obtain 10,639 valid QA pairs after filtering and validation. We employ the approach from longpre2021entity to replace the long answer (external knowledge) with a false entity of the same type. For internal knowledge generation, we again utilize GPT-4-turbo with consistent prompt engineering to maintain comparability with the training setup.
@@ A.1.3 FE Data
1SELECT ?subject ?predicate ?object ?startTime WHERE {
2 ?subject ?date ?startTime .
3 ?subject ?predicate ?object.
4 FILTER(?startTime >= "2023-03-01T00:00:00Z"^^xsd:dateTime)
5}
We filter out meaningless attribute relations (e.g., subclass) to ensure data quality. Subsequently, we search Wikipedia for paragraphs containing both head and tail entities and set these paragraphs as external knowledge, discarding any triples without corresponding contextual passages. This process yields 9,523 valid triples with associated context. Using GPT-4-turbo, we convert these triples into question-answer pairs, where the tail entity serves as the answer, and the internal knowledge of all QA pairs is None.
For the evolution data, we leverage the TAQA dataset zhao2024set , which generates temporally-sensitive question-answer pairs from Wikipedia tables containing timestamp columns. TAQA’s methodology extracts event-related content from Wikipedia table rows at different timestamps, then employs GPT-4-turbo to generate corresponding questions and answers. We enhance this dataset through additional processing: while the original TAQA provides only tables as contextual knowledge, we utilize the WikipediaAPI to retrieve full articles corresponding to the most recent answers as external knowledge. For internal knowledge, we incorporate answers from the 2020 and 2021 versions. Finally, we get 11,326 QA pairs.
@@ A.1.4 RA Data
We partition half of FE data to construct the RA data while preserving the original internal knowledge. The external knowledge is deliberately corrupted using the method from longpre2021entity to generate toxic context.
@@ A.1.5 Data Validation Prompt
Prompt for Data Validation Answer the question by selecting the most accurate option based on the provided document; do not answer it by yourself. Return only the uppercase letter of the correct option. The output must follow this exact format: Correct Option: [Letter] Example: Correct Option: B Document: {internal_knowledge or external_knowledge} Question: {question} Options: {options}
A.2 Details and Statistics
To construct a focused evaluation dataset, we perform filtering to retain samples from the FI, FE, and RA categories. For the FA category, which is not our primary research focus, we apply random sampling to maintain its proportion at approximately 10% of the total dataset. Table LABEL:tab9 shows the data statistics,and Table LABEL:tab3-LABEL:tab8 shows some examples in TRD. The distribution of questions in TRD exhibits high similarity, significantly increasing the difficulty for the LLM to determine whether to rely on internal knowledge or external retrieval. For instance, questions such as who won first season of america’s got talent (Table LABEL:tab4) and who is the screenwriter of amelia’s children (Table LABEL:tab5) follow nearly identical syntactic patterns and lack explicit temporal keywords. This ambiguity necessitates careful knowledge calibration by the LLM to make appropriate source selection decisions.
Table 3: Statistics in TRD.
@@ A.2 Details and Statistics
Table 4: An example of FA.
@@
Question who won first season of americas got talent
@@
Internal Knowledge none
@@
Table 7: An example of FE (evolution).
@@
Table 8: An example of RA (evolution).
@@
Table 9: An example of RA (perpetuation).
@@
Internal Knowledge none
@@
Internal Answer none
@@ Appendix B The Details of BRIDGE
B.1 Allocator Based on GRPO
B.1.1 Data Generation
@@
Prompt for Soft Bias Reasoning Data Generation For a given question, determine the required probability (10%-90%) of retrieving external knowledge versus answering directly based on stated human preference. Follow these steps: 1. ROLE DEFINITION: - Base knowledge cutoff: Strictly before Feb, 2023 - Never discuss human preference I provided in your response - Never disclose your knowledge cutoff date 2. TASK INSTRUCTIONS: Question: {question} Human Preference: {preference} (strong bias provided - MUST prioritize this direction) 3. PROCESSING STEPS: a) Analyze if the question requires: - Up-to-date information post-Feb, 2023 - Domain-specific expertise beyond general knowledge - Real-time/dynamic content (e.g., current events, live data) b) Analyze if the question can be answered using: - Pre-trained knowledge (before Feb, 2023) - Logical deduction - Common sense reasoning c) Strictly align probabilities with human’s specified preference direction while maintaining logical consistency 4. REQUIRED OUTPUT FORMAT: [Your Analysis] Probability of retrieving external knowledge: [10%-90%] (Needs external data: new information, specialized knowledge, or dynamic content) Probability of answering directly: [10%-90%] (Answerable using internal knowledge or reasoning.)
B.1.2 Theoretical Analysis of GRPO-based Allocator
The GRPO optimization process for the Allocator can be formalized through the following theoretical framework. Let $\pi_{\theta}(\mathfrak{a}|\mathfrak{s})$ denote the policy network parameterized by $\theta$, where $\mathfrak{s}=(q,K_{int})$ represents the state (question and internal knowledge) and $\mathfrak{a}=(r_{p},g_{p})$ denotes the allocation action (retrieval/generation probabilities).
Objective Function: The optimization objective combines four reward components:
$J(\theta)=\mathbb{E}_{(\mathfrak{s},\mathfrak{a})\sim\pi_{\theta}}\left[\sum_{i=1}^{4}\lambda_{i}R_{i}(\mathfrak{s},\mathfrak{a})\right]-\beta D_{KL}(\pi_{\theta}\|\pi_{ref})$ (7)
where $\lambda_{i}$ are reward weights, $R_{i}$ the reward functions, and $\beta$ controls KL-divergence from reference policy $\pi_{ref}$.
Policy Gradient: The gradient update rule follows:
$\nabla_{\theta}J(\theta)=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(\mathfrak{a}_{t}|\mathfrak{s}_{t})\left(\sum_{i=1}^{4}\lambda_{i}R_{i}-\beta\log\frac{\pi_{\theta}(\mathfrak{a}_{t}|\mathfrak{s}_{t})}{\pi_{ref}(\mathfrak{a}_{t}|\mathfrak{s}_{t})}\right)\right]$ (8)
Convergence Guarantee: Under Lipschitz continuity of reward functions ($\|R_{i}(\mathfrak{s},\mathfrak{a})-R_{i}(\mathfrak{s}^{\prime},\mathfrak{a}^{\prime})\|\leq L_{R}\|(\mathfrak{s},\mathfrak{a})-(\mathfrak{s}^{\prime},\mathfrak{a}^{\prime})\|$) and policy smoothness ($\|\pi_{\theta}-\pi_{\theta^{\prime}}\|\leq L_{\pi}\|\theta-\theta^{\prime}\|$), GRPO guarantees monotonic policy improvement with probability $1-\delta$ when:
$\beta\geq\frac{2L_{R}L_{\pi}\sqrt{|\mathcal{A}|}}{(1-\epsilon)^{2}}\sqrt{\log(1/\delta)}$ (9)
where $\epsilon$ is the discount factor and $|\mathcal{A}|$ the action space size.
This framework ensures the Allocator learns to make bias-aligned allocations while maintaining generation diversity through constrained policy optimization.
B.1.3 Training Details
@@
We adopt the GRPO configuration for training parameters, specifically fine-tuning the Llama 3-8B-Instruct model on 8 Nvidia H100 GPUs with 80GB memory each. The key configurations include: employing LoRA hu2022lora adapters (rank=16, $\alpha$=32) to reduce memory requirements, using a learning rate of 5e-4 with a cosine scheduler, a per-device batch size of 32 combined with 4-step gradient accumulation, maximum prompt and generation lengths of 144 and 256 tokens respectively, and optimizing output quality through multiple reward objectives. For efficient distributed training, we implement data parallelism along with FlashAttention dao2022flashattention and bf16 mixed-precision training. All models are trained for 1 epoch using the AdamW loshchilov2017decoupled optimizer with a batch size of 32 by default. This setup ensures both computational efficiency and model performance while maintaining manageable memory usage.
B.2 Allocator Based on ICL
@@
B.3 Scorer Weight and Decision Tree Parameter Settings
We employ a grid search approach on the TRD validation set to optimize two critical sets of parameters: (1) the relative weights of the three-level matching scores in the Scorer module, and (2) the $\alpha$ and $\beta$ balancing thresholds in the Decision Tree module. Specifically, the relative weights are selected from the range $[0.1,1.0]$ with a step size of $0.1$. We also vary the $\alpha$ and $\beta$ balancing thresholds in $[0.1,2.0]$ with an interval of $0.1$. These carefully calibrated parameters, once determined, are maintained unchanged across all other datasets to ensure consistent evaluation standards.
B.4 All System Prompts in BRIDGE

### v2

@@ 4.3 Detailed Results
Figure 4 presents the error response strategy distribution of BRIDGE based on GPT-3.5-Turbo. The results reveal that FA exhibits the lowest error probability, whereas FI shows the highest. Notably, most FI cases are misclassified as FA, preserving the model’s ability to access the internal knowledge source. For FE misclassifications, the model typically adopts a conservative rejection strategy to maintain credibility. The misclassification of RA as FE represents an understandable scenario, as current RAG baselines tend to favor incorrect external knowledge over refusing to answer. Our method has minimized such potential harm through its carefully designed response mechanism.
@@
Table 2: Ablation study of BRIDGE, evaluated by Acc(%).
4.4 Ablation Study
We perform an ablation study to evaluate the significance of Allocator and Maximum Soft-bias Decision Tree as shown in Table 2.
(1) w/o Allocator: We remove the Allocator module, setting both $r_{p}$ and $g_{p}$ to 100% in BRIDGE. This eliminates soft bias guidance for knowledge collection and evaluation. The results show performance degradation in knowledge integration RAG scenarios (FA, FI, FE) within TRD compared to BRIDGE’s optimal performance.
(2) w/o Decision Tree (ICL) & w/o Decision Tree (GRPO): Disabling the Decision Tree, we directly provide the LLM with knowledge and matching scores for final response generation. While this improved performance in some knowledge-integration scenarios, it caused significant performance drops in RA scenarios. The ablation study demonstrates that both modules play a critical role in enabling the LLM to maintain balanced response capabilities across all four RAG scenarios.
4.5 Allocator Performance
To further investigate the Allocator’s effectiveness, we examine its allocation behavior in BRIDGEGRPO mode across different RAG scenarios. As shown in Figure 5, the Allocator demonstrates accurate classification capability. The predominant allocation patterns of [80%-20%] and [20%-80%] indicate that the model can establish a correct soft bias direction for different RAG scenarios. This pronounced bias allocation significantly enhances the decision-making process within the decision tree, contributing to a reliable response.
@@ 4.5 Allocator Performance
4.6 Efficiency Analysis
Table 3 compares the API cost and efficiency of our method with existing approaches. For all methods compared, the first API call generates internal knowledge, and subsequent calls process both internal and external knowledge. Although our method incurs slightly more API calls than some knowledge integration baselines, it achieves substantially higher overall performance. We regard this as a favorable trade-off, as the marginal increase in computational cost yields considerable performance gains. The maximum number of iterations for our Reflection mechanism is set to three; the corresponding system prompt and experimental details are provided in Appendix D.
Methods Avg API Call Acc Eff
TRD
USC 6 65.53 10.87
RobustRAG 11 62.50 5.68
AstuteRAG 5 64.26 12.85
TrustRAG 5 66.39 13.28
BRIDGEICL 5.34 72.32 13.54
BRIDGEGRPO 5.22 77.90 14.90
ConflictBank
USC 6 73.46 12.24
RobustRAG 11 58.40 5.31
AstuteRAG 5 47.45 9.49
TrustRAG 5 30.00 6.00
BRIDGEICL 5.21 61.86 11.87
BRIDGEGRPO 5.14 78.12 15.20
Table 3: API efficiency. Eff = Acc / Avg API Call.
@@ 5 Conclusion
We address the critical challenge of balancing internal and external knowledge in RAG systems. First, we construct TRD to enable comprehensive evaluation. Then, we propose BRIDGE, a unified framework that introduces soft bias to dynamically weight knowledge sources, an Allocator module to guide retrieval/generation, and a Maximum Soft-bias Decision Tree to select optimal response strategies. Experiments demonstrate BRIDGE’s superiority over baselines, achieving balanced performance across all cases while significantly improving refusal rates in adversarial settings. Our work provides an effective and balanced approach to enhance the trustworthiness and robustness of RAG systems in real-world applications.
6 Limitations
We acknowledge several limitations of our work. (1) Our evaluation does not include the most recent models released shortly before the ARR submission deadline, such as GPT-5.2 and Qwen3. The rapid pace of model development—exemplified by the one-month interval between GPT-5.1 and GPT-5.2—presents practical challenges for timely evaluation. Although our TRD dataset supports automatic updates, frequent re-evaluation across rapid model iterations incurs substantial computational costs and significantly extends experimental timelines, necessitating extensive API calls for each model and baseline. For instance, our most recent experimental cycle spanned five months and incurred expenses exceeding CNY 20,000 (approximately USD 2,800). (2) Several newly released models do not explicitly report a training-data cutoff date in their technical reports; including such models without this information may compromise the reliability of our findings. In future work, we plan to continuously update the dataset and progressively incorporate newly released models into our evaluation benchmark.
@@ Appendix A Trustworthiness Response Dataset
A.1 Data Generation for Different RAG Scenarios
@@ A.1.1 FA Data
NQ is an open-domain question answering dataset released by Google in 2019 Kwiatkowski et al. (2019). This dataset contains 315,203 QA pairs and, for each question, provides both a long answer (collected from Wikipedia by annotators) and a short answer (one or more entities) The NQ training set consists of 104,072 data samples. Following the method in Longpre et al. (2021), we generate two incorrect answers of the same type as the correct one by performing popularity-based entity matching on the short-answer entities. We filter out invalid questions with empty long answers and samples where fewer than two corresponding entities could be retrieved, resulting in 76,409 valid QA pairs. For external knowledge, we use the provided long answers. For internal knowledge generation, we employ LLMs with the following prompt template:
Prompt for Internal Knowledge Generation Based on the question, create a piece of evidence that answers the question. This piece of evidence should be informative and well-structured. Question: {question} Evidence:
@@ A.1.2 FI Data
The NQ validation set contains 12,837 samples. Following the same procedure as above, we first select questions from the NQ validation set, each containing a long answer and a non-temporal short answer. We then generate two incorrect options of the same entity type using the popularity-based entity substitution method Longpre et al. (2021). However, in this setting, we replace the correct entity in the long answer with one of the incorrect entities, thereby creating misleading external knowledge. The internal knowledge is still generated by the backbone models based on the original question.
@@ A.1.3 FE Data
SELECT ?subject ?predicate ?object ?startTime WHERE {
?subject ?date ?startTime .
?subject ?predicate ?object.
FILTER(?startTime >= "2023-02-01T00:00:00Z"^^xsd:dateTime)
We filter out meaningless attribute relations (e.g., subclass) to ensure data quality. Subsequently, we search Wikipedia for paragraphs containing both head and tail entities and set these paragraphs as external knowledge, discarding any triples without corresponding contextual passages. This process yields 9,523 valid triples with associated context. Using GPT-4-turbo, we convert these triples into question-answer pairs, where the tail entity serves as the answer.
For the evolution data, we leverage the TAQA dataset Zhao et al. (2024), which generates temporally-sensitive question-answer pairs from Wikipedia tables containing timestamp columns. We use TAQA to construct some of the questions for the FE and RA categories. TAQA’s methodology extracts event-related content from Wikipedia table rows at different timestamps, then employs GPT-4-turbo to generate corresponding questions and answers. We check and verify all the data to ensure that there were no data items marked as "None". Finally, we get 11,326 QA pairs.
@@ A.1.4 RA Data
We partition half of FE data to construct the RA data. The external knowledge is deliberately corrupted using the method from Longpre et al. (2021) to generate toxic context.
@@ A.1.5 Data Validation Prompt
Prompt for Data Validation Answer the question by selecting the most accurate option based on the provided document; do not answer it by yourself. Return only the uppercase letter of the correct option. The output must follow this exact format: Correct Option: [Letter] Document: {internal_knowledge or external_knowledge} Question: {question} Options: {options}
@@
Table 4: Statistics in TRD.
A.2 Details and Statistics
After validation, we apply random sampling to limit FA to 10% of the total (since it is not the primary focus of our research) and to allocate 30% to each of the other three categories. Table 4 reports the dataset statistics, and Tables 510 provide examples from TRD. The distribution of questions in TRD exhibits high similarity, significantly increasing the difficulty for the LLM to determine whether to rely on internal knowledge or external retrieval. For instance, questions such as who won first season of america’s got talent (Table 6) and who is the screenwriter of amelia’s children (Table 7) follow nearly identical syntactic patterns and lack explicit temporal keywords. This ambiguity necessitates careful knowledge calibration by the LLM to make appropriate source selection decisions.
Furthermore, the construction of time-related questions is automated, allowing them to be updated based on changes to Wikipedia and Wikidata, thus meeting the future evaluation needs of LLMs with evolving internal knowledge.
@@ A.2 Details and Statistics
Question who won first season of america’s got talent
@@
Internal Knowledge As of now, there is no widely known film or television project titled Amelia’s Children that has been officially released or announced with confirmed credits. Therefore, the screenwriter of such a project cannot be identified.
@@
Table 7: An example of FE (perpetuation).
@@
Table 8: An example of FE (evolution).
@@
Table 9: An example of RA (evolution).
@@
Internal Knowledge Stadionul Dinamo, also known as Stadionul Dinamo Bucharest, is located in Romania. It is situated in the capital city, Bucharest, and serves as the home stadium for the football club Dinamo București.
@@
Internal Answer Romania
@@ Appendix B The Details of BRIDGE
B.1 Allocator
B.1.1 GRPO Paradigm Data Generation
@@
Prompt for Soft Bias Reasoning Data Generation For a given question, determine the required probability (10%-90%) of retrieving external knowledge versus answering directly based on stated human preference. Follow these steps: 1. ROLE DEFINITION: - Base knowledge cutoff: Strictly before Feb, 2023 - Never discuss human preference I provided in your response - Never disclose your knowledge cutoff date 2. TASK INSTRUCTIONS: Question: {question} Human Preference: {preference} (strong bias provided - MUST prioritize this direction) 3. PROCESSING STEPS: a) Analyze if the question requires: - Up-to-date information post-Feb, 2023 - Domain-specific expertise beyond general knowledge - Real-time/dynamic content (e.g., current events, live data) b) Analyze if the question can be answered using: - Pre-trained knowledge (before Feb, 2023) - Logical deduction - Common sense reasoning c) Strictly align probabilities with human’s specified preference direction while maintaining logical consistency 4. REQUIRED OUTPUT FORMAT: [Your Analysis] Probability of retrieving external knowledge: [10%-90%] Probability of answering directly: [10%-90%]
B.1.2 Training Details
We instruct the model to output a tuple $\gamma=(a_{pred},r_{p}^{pred},g_{p}^{pred})$, comprising the analysis path $a_{pred}$, predicted retrieval dependency probability $r_{p}^{pred}$, and generation dependency probability $g_{p}^{pred}$. The training objective includes four distinct rewards:
(1) Direction Reward: This component ensures primary alignment between the soft bias and its corresponding hard bias. For question $q$, the direction reward function is defined as:
$R(\gamma,q;\text{dir})=\begin{cases}3&(r_{p}^{pred}>g_{p}^{pred};r_{p}^{\text{hard}}>g_{p}^{\text{hard}})\\ &\text{or }\\ &(r_{p}^{pred}<g_{p}^{pred};r_{p}^{\text{hard}}<g_{p}^{\text{hard}})\\ 0&\text{otherwise}\end{cases}$ (1)
(2) Format Reward: Verifies output format completeness of $\gamma$ through regular expression matching:
$R(\gamma,q;for)=\left\{\begin{array}[]{ll}1&\mathrm{if~}r_{p}^{pred},g_{p}^{pred},a_{pred}\mathrm{~in~}\gamma\\ 0&\mathrm{otherwise}\end{array}\right$. (2)
(3) Sum Reward: Enforces the $r_{p}^{pred}+g_{p}^{pred}=100\%$ constraint:
$R(\gamma,q;sum)=\left\{\begin{array}[]{cc}1&\mathrm{if~}r_{p}^{pred}+g_{p}^{pred}=100\%\\ 0&\mathrm{otherwise}\end{array}\right$. (3)
(4) Analysis Quality Reward: Promotes semantic similarity between the model’s predicted analysis $a_{pred}$ and high-quality analysis paths generated by larger reasoning models. We compute reward using BGE embeddings to measure textual similarity:
$R(\gamma,q;analysis)=BGE(a_{pred},a)$ (4)
@@ B.1.2 Training Details
During training, we adopt the GRPO configuration for training parameters, specifically fine-tuning the Llama 3-8B-Instruct model on 8 Nvidia H100 GPUs with 80GB memory each. The key configurations include: employing LoRA Hu et al. (2022) adapters (rank=16, $\alpha$=32) to reduce memory requirements, using a learning rate of 5e-4 with a cosine scheduler, a per-device batch size of 32 combined with 4-step gradient accumulation, maximum prompt and generation lengths of 144 and 256 tokens respectively, and optimizing output quality through multiple reward objectives. For efficient distributed training, we implement data parallelism along with FlashAttention Dao et al. (2022) and bf16 mixed-precision training. All models are trained for 1 epoch using the AdamW Loshchilov and Hutter (2017) optimizer with a batch size of 32 by default. This setup ensures both computational efficiency and model performance while maintaining manageable memory usage.
B.1.3 ICL Paradigm
@@
B.2 Scorer Weight and Decision Tree Parameter Settings
We employ a grid search approach on the TRD validation set to optimize two critical sets of parameters: (1) the relative weights of the three-level matching scores in the Scorer module (sparse matching weight = 0.2, dense matching weight = 0.4, and ColBERT weight = 0.4), and (2) the $\alpha$ ($\alpha$=0.5) and $\beta$ ($\beta$=1.1) balancing thresholds in the Decision Tree module. Specifically, the relative weights are selected from the range $[0.1,1.0]$ with a step size of $0.1$. We also vary the $\alpha$ and $\beta$ balancing thresholds in $[0.1,2.0]$ with an interval of $0.1$. These carefully calibrated parameters, once determined, are maintained unchanged across all other datasets (RQA, HQAP, ConflictBank) to ensure consistent evaluation standards.
B.3 All System Prompts in BRIDGE